OpenStudy (anonymous):

How to find a vector that is perpendicular to another? I know their dot product = 0, but is there was a way of finding a perpendicular vector right away?

OpenStudy (anonymous):

to check for perpendicular dot product is the best one

OpenStudy (anonymous):

yes, but say for example i have <2/3, 1/2, 2/3> how do i find a perpendicular vector? I mean, I'm guessing numbers atm which I know is wrong. Is there a better way?

OpenStudy (anonymous):

can u state ur complete question please

OpenStudy (anonymous):

Ok. What is a vector that is perpendicular to the following unit vector <2/3, 1/2, 2/3>?

OpenStudy (anonymous):

you can make the dot product equal zero in this case you have x=2/3 and z=2/3 set one to 1 and the other to -1 and y=0 <-1,0,1> will give you a perpendicular vector I don't know if there is a better way